The area of a rectangular wall of a barn is 126 feet.Its length is 4 feet longer than twice its width. Find the length and width of the wall of the barn
Answer by math_helper(2461)

The area of a rectangular wall of a barn is 126 feet.Its length is 4 feet longer than twice its width. Find the length and width of the wall of the barn
===================================
126 ft is not a valid representation of area. 126 square feet (sq ft or ft^2) maybe.
Area = Length * Width
126 sq ft = (2W+4)(W)
2W^2 + 4W - 126 = 0
W^2 + 2W - 63 = 0
(W-7)(W+9) = 0

W = 7 or W = -9 (discard -9 as Width must be > 0)
W = 7 —> L = 2(7)+4 = 18
===
Ans: Width is 7ft, Length is 18ft
===
Check (always a good idea, especially for me): 7ft*18ft = 126ft^2 (ok)
Also, 18 = 2*7 + 4 = 14+4 (ok)